### v1

@@ 5.2 Hallucinating a stopped vehicle ahead
At another location, the AV would reliably come to a stop next to a traffic cone (Figure 3b). The driver’s initial mental model was that the cone was responsible for the phantom brake. However, the “Approaching stopped vehicle” (ASV) concept peaked shortly before the car stopped. This suggested an alternative hypothesis, that the planner matched the current situation with training scenarios labeled ASV, which in turn promotes stopping behavior associated with such scenarios. As a counterfactual test, the cone was removed. The AV exhibited the same stopping behavior at the same location, along with similar ASV probability and speed profiles ($L_{2}$ similarities of 1.12 and 1.6 between the respective time-warped profiles, compared to an average $L_{2}>200$ for random events), thus confirming the alternative hypothesis. Note that although there was no vehicle in front of the AV, the explanation is causally faithful to the underlying motion planner and explains why it stopped (namely, because it incorrectly detected a stopped vehicle). Figure 3b illustrates a global analysis of ASV, showing it to be a powerful predictor of braking.
@@ 5.3 Reacting safely to cyclist
Finally, we tested the ability of the AV to stop safely for cyclists (Figure 3c). For each test, the driver engaged self-driving mode while approaching a cyclist. The driver was instructed to engage self-driving from a speed at which they felt safe, since this determines the subsequent speed of the AV. During the initial tests, the AV reliably stopped for the cyclist. However, the BIKE concept maintained a low probability throughout each test ($<$ 1%), indicating that CW-Net was failing to detect the cyclist. Over time, the driver became aware of the concept reading and gradually increased their caution by initiating self-driving from slower speeds. A post-hoc analysis revealed that, although the perception system detected the cyclist, the motion planner failed to take it into account due to a lack of appropriate input features for cyclists. As a result, it chose unsafe trajectories which would have collided with the cyclist. In reality, it was the built-in rule-based systems of the AV that overruled the motion planner and forced the AV to stop. This indicates that the increased caution dictated by the driver’s updated mental model was warranted.
6 Confirmation study
We conducted an online study (N=120) to confirm the statistical significance of our results. Following standard procedures in deployment-based research [FAA2011FlightTest, schneider2023don], we simulated the sequence of key events from the three situations described previously. We designed a between-subjects study to assess the effect of explanations on the mental model of subjects in the driver’s position. For each situation, subjects in the experimental group received the CW-Net explanation, while subjects in the control group received a generic explanation to balance cognitive load (see Figure S4 and Section User Study for details). Subjects then rated to what extent they agreed with the driver’s incorrect initial mental model. Across all situations, the experimental group gave significantly lower ratings than the control group (Figure 3, right column; mean result: 3.37$\pm$1.63 v. 5.46$\pm$0.89; t-test $p$ $<$ 0.001; Cohen’s d=1.58), indicating that the explanations support mental model alignment across the population.
@@ 7 Conclusion
Our work shows, for the first time, how explainable deep learning can provide useful explanations for the decisions of self-driving cars in the real world. CW-Net achieves this by grounding the reasoning of a pretrained black-box motion planner in human-interpretable concepts corresponding to types of scenario. By revealing otherwise inaccessible information about the decision-making process of the motion planner in real time, CW-Net helps align the mental model of the human driver with the machine driver. This allows the human driver to better anticipate and account for mistakes of the AV, ultimately resulting in safer driving. Mental model alignment could additionally build trust and understanding with passengers of driverless AVs by helping them anticipate the AV’s decisions. This could be particularly beneficial when AV behavior deviates from typical human behavior, such as when driving conservatively or getting stuck. Additionally, the explanations provided by CW-Net can help test engineers provide more precise feedback to the research scientists and engineers working on model improvement. This would be especially relevant for experimental motion planners that are still under development, such as the one used in our study.
CW-Net extends the original PW-Net [kenny2023towards], which reasons over specific scenario prototypes, to general scenario types. In addition to increasing robustness, reasoning over types has the added benefit of highlighting which parts of the training distribution influence behavior at each point in time. This information can be used by researchers for model improvement. For example, the inability of CW-Net to detect the BIKE concept suggests there may not be enough training scenarios with cyclists, or that the focal loss used was ineffective [lin2017focal], leading to poor performance around cyclists. At the same time, relying on types forgoes some of the benefits of using prototypes. For example, CW-Net can explain that it is stopping because the current scene is similar to ASV scenarios in the training data, but it cannot explain why it believes so  [rudin2019stop, chen2019looks]. This suggests a promising avenue of future research: much like humans, who rely on both exemplar-based and rule-based reasoning [smith1994similarity, davis2012learning], machines that reason over both prototypes and types could combine the strengths of both approaches, enhancing interpretability while maintaining flexibility in decision making.
Many safety-critical systems involving human-robot interaction require real-time explanations, including AI wingmen, drone navigation systems, and robotic surgeons. Similarly to AVs, many of these applications increasingly rely on deep learning and have a long tail of failure cases, with potentially catastrophic outcomes. This creates an ethical imperative to better understand how they work, so the humans in the loop can intervene when necessary. The success of CW-Net suggests that explainable deep learning may prove essential for meeting the safety and regulatory standards for deploying sophisticated safety-critical agents in the real world.
@@ Acknowledgements
Figure S1: Parallel architecture: This is identical to the black-box planner (Figure 1b), except the scene-trajectory embeddings are fed to the concept classifier $C$ in parallel to the (original) reward model $R$. These concept classifications are then converted to probabilities (x100 to convert to percentages) and presented to the user.
Scenario Metric Our Model Original
nominal lane follow (968) Average L2 error 4.202058 4.200189
@@
decelerating from high speed scenarios (1099) Average L2 error 3.139028 3.144618
@@
start accelerating from stationary scenarios (919) Average L2 error 8.919248 9.019697
@@
Table S1: Full nuPlan results which compares the performance of our main model used in testing against the original black-box IRL agent.
@@
Approaching stopped vechicle 0.55 0.2 0.89 0.33
@@
CLOSE to another vechicle 0.45 0.1 0.81 0.17
@@
Table S2: Concept performance for Dataset 1 for CW-Net
@@
Table S3: Concept performance for dataset 1 with CW-Net.
Figure S2: A comparison of the concept predictions for the same model deployed in a large test track with wide open, long straight roads, and a smaller parking lot with many turns.
Figure S3: Statistical results for user study: (left) Each individual question.(right). Averaged across all questions
Figure S4: Survey questions for the experimental and control group.

### v2

@@ 5.2 Hallucinating a stopped vehicle ahead
At another location, the AV would reliably come to a stop next to a traffic cone (Figure 3b). The driver’s initial mental model was that the cone was responsible for the phantom brake. However, the “Approaching stopped vehicle” (ASV) concept peaked shortly before the car stopped. This suggested an alternative hypothesis, that the planner matched the current situation with training scenarios labeled ASV, which in turn promotes stopping behavior associated with such scenarios. As a counterfactual test, the cone was removed. The AV exhibited the same stopping behavior at the same location, along with similar ASV probability and speed profiles ($L_{2}$ similarities of 1.12 and 1.6 between the respective time-warped profiles, compared to an average $L_{2}>200$ for random events), thus supporting the alternative hypothesis. Note that although there was no vehicle in front of the AV, the explanation is causally faithful to the underlying planner and explains why it stopped (namely, because it incorrectly detected a stopped vehicle). Figure 3b illustrates a global analysis of ASV, showing it to be a powerful predictor of braking.
@@ 5.3 Reacting safely to cyclist
Finally, we tested the ability of the AV to stop safely for cyclists (Figure 3c). For each test, the driver engaged self-driving mode while approaching a cyclist. The driver was instructed to engage self-driving from a speed at which they felt safe, since this determines the subsequent speed of the AV. During the initial tests, the AV reliably stopped for the cyclist. However, the BIKE concept maintained a low probability throughout each test (<< 1%), indicating that CW-Net was failing to detect the cyclist. Over time, the driver became aware of the concept reading and gradually increased their caution by initiating self-driving from slower speeds. A post-hoc analysis revealed that, although the perception system detected the cyclist, the ML planner was not configured to consume inputs for cyclists. As a result, it chose unsafe trajectories which would have collided with the cyclist. In reality, the AV stopped due to a built-in safety backup system, which commands a brake if collision is imminent. This indicates that the increased caution dictated by the driver’s updated mental model was warranted.
6 Mental model improvement in simulation studies
The semi-naturalistic study above illustrates how CW-Net explanations can improve the human driver’s mental model of the AV in surprising real-world situations, leading to better understanding and predictability of AV behavior. To validate that these results replicate in larger populations and generalize to naturalistic scenarios, we conduct several larger follow-up studies (Figures S2S3). First, we show that in simulations of the real-world scenarios in Section 5, CW-Net explanations consistently improve mental models and predictions for both experts and non-experts alike. Second, we collect data from CW-Net in naturalistic real-world scenarios on public roads and show that the explanations improve situational awareness – an indirect measure of mental model goodness [12] – in surprising situations, without degrading it in unsurprising situations.
6.1 Explanations improve mental models and predictability
The purpose of the first study is to (1) demonstrate that the effects observed on the road reproduce in larger populations, and (2) establish a link between direct measures of mental model goodness and performance on downstream tasks, such as predicting AV behavior. We conducted an online survey in which we showed participants replays of the scenarios from the private track (front camera recordings with overlaid CW-Net explanations; Figure S11).
For each scenario, we asked participants to choose the reason for AV behavior that best aligns with their current beliefs (i.e., “Why did the AV do that?”) and to make counterfactual predictions (i.e., “What would the AV do if …?”). Following Hoffman et al. [11], we refer to the former as the nearest-neighbor task – since it forces participants to pick the nearest explanation to their beliefs – and the latter as the prediction task. Both tasks consist of a multiple-choice question followed by a confidence score. The prediction task additionally includes a free-form text response where participants can describe the reasons for their prediction. The nearest-neighbor task and the free-form text response directly probe the participants’ mental models, while the multiple-choice prediction question serves as a proxy performance metric for mental model goodness. Critically, we collected responses before and after observing the explanations. This within-subject design mirrors the experience of drivers during the real-world tests (Figures S2, S3). The study was conducted with experts ($N=9$ Motional drivers/engineers) and non-experts ($N=30$ pseudo-random participants from Prolific.com).
On the nearest-neighbor task, CW-Net explanations improved mental models for almost all participants (8/9 experts and 27/30 non-experts; Figure S4). Free-form text responses were initially similar to the initial beliefs of the safety drivers during the on-road tests ($p<10^{-9}$) and then shifted towards their final beliefs ($p<0.0002$) and the ground-truth reasons for AV behavior ($p<10^{-5}$; exact binomial tests for non-experts; Figures S6S7). The distribution of mental model updates on both tasks was similar across both groups ($\beta=0.04,p=0.8$, ordinary least-squares regression; Figure S8). Importantly, mental model improvement on the nearest-neighbor task correlated with improvements in prediction accuracy ($\beta=2.02\pm 0.87,p=0.02$ for experts, $\beta=9.86\pm 2.07,p<0.001$ for non-experts; linear mixed-effects models [LMEs]; Figure S9, left). A similar effect was observed for the free-form text responses ($\beta=1.70\pm 0.91,p=0.06$ for experts, $\beta=5.03\pm 1.23,p<0.001$ for non-experts; LMEs; Figure S9, right). Together, these results support the conclusion that CW-Net explanations can reliably improve mental models of the AV and that prediction performance can serve as a reliable proxy for mental model goodness.
6.2 Explanations improve situational awareness in naturalistic scenarios
To verify that the effects observed on the private track generalize to more complex, naturalistic scenarios, we deployed CW-Net on public roads in Las Vegas (Figures S3; Figure 3, right). Due to safety reasons and the experimental nature of the ML planner, the safety driver operated the AV in ‘manual’ mode for several hours, with CW-Net running in the background. This allowed us to collect naturalistic scenarios analogous to those encountered on the private track [41, 28] (Figure S10). We use replays of those scenarios to conduct a large-scale online study ($N=100$) within the Situation Awareness Global Assessment Technique (SAGAT) framework tailored for explainable AI [12, 42].
The SAGAT framework is the gold standard for evaluating situational awareness in complex, dynamic tasks [12, 42]. It measures the situational awareness of a human operator by freezing a scenario, querying the operator about their perception (i.e., “What are the inputs to the AV?”), comprehension (i.e., “Why is the AV doing that?”), and projection (i.e., “What would the AV do if …?”) of the situation (Figure S11) and comparing their responses with the ground truth. As counterfactual predictive performance in Section 6.1 aligns with mental model goodness, we use projection in SAGAT as a proxy for mental model elicitation and quality [11]. For each type of scenario and corresponding concept, we sample two instances in which AV behavior was surprising (e.g., stopping unnecessarily) and two corresponding instances in which it was unsurprising (Table S6). We use a between-subject design in which the experimental group received explanations from CW-Net, while the control group received placeholder explanations of AV kinematics to balance cognitive load [14] (Figures S12-S14).
We found that the CW-Net explanations significantly improved measures of situational awareness for surprising events (Figure S15, left; Table S7, top), with large effect sizes for perception (Cohen’s d = 1.290), comprehension (Cohen’s d = 0.996), and a medium effect size for projection (Cohen’s d = 0.606). Conversely, explanations did not impact situational awareness for unsurprising events (Figure S15, right; Table S7, bottom). Together, these results show that the explanations from CW-Net are robust across various conditions and can reliably improve mental models in surprising situations, without significant negative effects in unsurprising situations. Moreover, they demonstrate pragmatic usage of the improved mental models.
@@ 7 Conclusion
Our work shows, for the first time, how explainable deep learning can provide useful explanations for AVs in a real-world setting. CW-Net achieves this by grounding the reasoning of a pretrained black-box ML planner in human-interpretable concepts that are directly used to make driving decisions. By revealing otherwise inaccessible information about the decision-making process of the AV in real time, CW-Net helps improve the human driver’s mental model of the AV. This, in turn, improves the driver’s situational awareness and reveals limitations of the robotic system, helping the driver better anticipate its mistakes.
Many systems involving human-robot interaction require real-time explanations, and many of these applications increasingly rely on deep learning, with a long tail of potentially catastrophic failure cases. Indeed, many regulatory bodies have already made explainable AI a core component of their legislation, with AVs likely to follow suit as they are widely deployed with various users [11]. As such, the success of CW-Net suggests that similar algorithms may prove essential for meeting the regulatory standards for deploying AVs, while building appropriate trust in the technology. In future work, it would be prudent to extend CW-Net to cover a larger set of concepts – perhaps in an unsupervised manner to overcome the challenges of labeling – to better cover the vast array of concepts relevant to AV settings.
@@ Acknowledgements
Figure S1: Parallel architecture: This is identical to the black-box planner (Figure 1b; greyed out here), except the scene-trajectory embeddings are fed to the concept classifier $C$ in parallel to the (original) reward model $R$. The black-box part of the architecture (greyed out) is kept frozen. These concept classifications are then converted to probabilities (x100 to convert to percentages) and presented to the user.
Scenario type Metric CW-Net ML planner
Nominal lane follow (968 scenarios) Average L2 error 4.202058 4.200189
@@
Decelerating from high speed (1099 scenarios) Average L2 error 3.139028 3.144618
@@
Start accelerating from stationary (919 scenarios) Average L2 error 8.919248 9.019697
@@
Table S1: Full nuPlan results comparing driving performance of our CW-Net planner to the original black-box ML planner.
@@
Approaching stopped vechicle (ASV) 0.55 0.20 0.89 0.33
@@
CLOSE to another vechicle 0.45 0.10 0.81 0.17
@@
Table S2: CW-Net concept classification performance on Dataset 1
@@
Table S3: CW-Net concept classification performance on Dataset 2.
Figure S2: Study methodology. First we collected a series of surprising situations expert safety drivers encountered during their day-to-day job testing the AV where the explanations proved to be useful to understand why the AV made certain decisions and how it would behave in alternative scenarios. During these situations, we noted the driver’s mental model through their think-aloud thoughts and behaviour. A mental model elicitation study was run with more expert drivers and non-experts online to more robustly verify these findings and if mental model goodness did indeed correlate with counterfactual predictive performance. Then, in a large-scale SAGAT study, we used projection as a surrogate metric for mental model quality, since it perfectly matches the prior prediction tasks.
Figure S3: Study pipeline example. We noted the safety drivers’ think-aloud thoughts and predictions in surprising situations encountered during the on-road tests. In the subsequent online mental model elicitation study, we used their beliefs about AV behaviour before/after the explanation in the nearest-neighbor task and their predictions about future AV future behavior in the prediction task. The nearest-neighbor task resembles the comprehension task in the follow-up SAGAT study. The prediction task corresponds exactly to the projection task in the SAGAT study.
Scenario/concept Before explanation After explanation
CLOSE 0.00% 11.11%
ASV 33.33% 44.44%
BIKE 77.78% 77.78%
Overall average 37.04% 44.44%
Table S4: Performance improvement on prediction task ($N=9$ experts).
Scenario/concept Before explanation After explanation
CLOSE 3.33% 33.33%
ASV 36.67% 60.00%
BIKE 40.00% 66.67%
Overall average 26.67% 53.33%
Table S5: Performance improvement on prediction task ($N=30$ non-experts).
Figure S4: Mental model improvement on nearest-neighbor task. Visualizing belief updates in response to the explanations for experts (left) and non-experts (right) on the nearest-neighbor task. Arrows originate at the participants’ initial accuracy/confidence (before observing the explanation) and terminate at their final accuracy/confidence (after observing the explanation). For each participant, initial and final accuracy/confidence is averaged across all 3 scenarios. Green arrows indicate improvement in choice accuracy/confidence, while red arrows indicate worsening. Gray points represent participants with no change in beliefs.
Figure S5: Performance improvement on prediction task. Visualizing belief updates in response to the explanations for experts (left) and non-experts (right) on the prediction task. Notation as in Figure S4.
Figure S6: Mental model improvement on free-form text rationale ($N=9$ experts). Confusion matrices with participant beliefs (mental model and predictions) extracted from free-form responses in the mental model elicitation study study (y-axis) compared to beliefs from the safety drivers from the on-road tests and the ground truth explanation for AV behavior (x-axis). Numbers indicate raw response counts.
Figure S7: Mental model improvement on free-form text rationale ($N=30$ non-experts). Notation as in Figure S6.
Figure S8: Correlation between mental model metrics across groups. Heatmaps displaying the frequency of alignment between changes in nearest-neighbor (NN) task responses and the free-form text rationale. To determine if the relationship between these two mental model proxies differs by expertise, we performed ordinary least-squares (OLS) regression with clustered standard errors, including an interaction term. The interaction between group (experts vs. non-experts) and score change was not significant ($\beta=0.04,p=0.842$), indicating that the relationship between quantitative score improvements and qualitative rationale improvements is largely consistent across both groups.
Figure S9: Mental model improvement correlates with prediction improvement. Linear mixed effects (LME) models were used to analyze the relationship between changes in mental model proxies (nearest-neighbor task and free-form text rationale) and prediction accuracy/confidence. Specifically, the y-axis indicates each participant’s Likert confidence after explanation minus before, positive scores indicate they moved towards the correct prediction, and green dots indicate the prediction flipped correctly. For experts ($N=9$ participants x 3 scenarios), improvement in the nearest-neighbor (NN) task significantly correlates with prediction improvement ($\beta=2.02,SE=0.87,p=0.02$). The text rationale shows a similar positive effect ($\beta=1.70,SE=0.91,p=0.06$), although it does not reach significance. For non-experts ($N=30$ participants x 3 scenarios), both NN improvement ($\beta=9.86,SE=2.07,p<0.001$) and text rationale improvement ($\beta=5.03,SE=1.23,p<0.001$) are significantly correlated with prediction improvement. Positive betas indicate that improving the mental model proxy is associated with a higher increase in prediction accuracy/confidence.
Figure S10: Naturalistic scenario examples. The original surprising scenarios discovered on the private track and the analogous scenarios discovered on public roads.
Figure S11: SAGAT materials example. The two conditions (explanation vs. no explanation) were presented with the same video, followed by a “blackout”, and then 6 questions related to situational awareness.
Testing concept robustness
Concept activation AV speed Concept accuracy
Original
ASV High Brakes Bad
CLOSE High Frozen Good
BIKE Low Moving Low
Modified for robustness
ASV High Brakes Good
CLOSE Low Frozen Good
BIKE Low Stopped Low
Table S6: Selecting unsurprising events for SAGAT study. To select unsurprising events we modifed what was individual unusual about each surprising event which fell into three categories. ASV: This concept was surprising because of the misclassification of the concept, so we selected a scenario in which it accurately classified the concept. CLOSE: This was surprising because the concept activation was higher than usual while the AV stopped, so we chose scenarios in which the AV stopped, but the concept activation was closer to its mean value. BIKE: This was unusual because the AV stopped seemingly due to the cyclist, so we chose scenarios in which the AV again stopped for other reasons. In all three setups, we only altered the one respective dimension, the other two remained static.
Figure S12: User materials example for the CLOSE concept, during a surprising and unsurprising event.
Figure S13: User materials example for the ASV concept, during a surprising and unsurprising event.
Figure S14: User materials example for the PEDESTRIAN concept, during a surprising and unsurprising event.
Figure S15: Online SAGAT study results. Users in the experimental group reported significantly higher accuracy on perception, comprehension, and projection during surprising events. During unsurprising events, there were no significant differences. Although a negative trend was noted for comprehension in the experimental group, it did not reach significance after Bonferroni correction (see Table S7). Taken with the results of the mental model elicitation studies, which found counterfactual prediction accuracy correlated with mental model goodness, these SAGAT results strongly indicate the improvements in situational awareness were due to underlying improvements in user mental models of the AV.
Valence Dimension Effect Cohen’s d $p$-value Bonferroni
direction significant
Surprising events Perception Exp >> Control 1.290 (large) <0.001<0.001 Yes
Comprehension Exp >> Control 0.996 (large) <0.001<0.001 Yes
Projection Exp >> Control 0.606 (medium) 0.003 Yes
Unsurprising events Perception No effect 0.085 (negligible) 0.667 No
Comprehension Control >> Exp 0.514-0.514 (medium) 0.011 No
Projection No effect 0.142-0.142 (negligible) 0.481 No
Table S7: SAGAT study results. During the ‘surprising’ events, explanations had an all-around positive impact on people’s situational awareness and thus their mental models. Conversely, during ‘unsurprising’ events, there were no significant differences between groups, after Bonferroni correction.
FEATURE Model 1 Model 2 Model 3
STOPPED 0.0862 0.0014 0.0324
SLOW 0.1575 0.0014 0.0324
FAST 0.1903
STOP SIGN 0.2368
TRAFFIC LIGHT 0.0938
INTERSECTION 0.0856 0.0247 0.0535
PEDESTRIAN 0.0998
FOLLOWING 0.1056
BIKE 0.1222
PUDO 0.2347
LEFT 0.0472 0.0332
RIGHT 0.0408 0.0410
STRAIGHT 0.0428 0.0592
ASV 0.0269 0.1050
CLOSE 0.0659 0.0702
Table S8: Negligible concept distribution shift: Wasserstein distances between the distribution of concept probabilities during the semi-naturalistic tests on the private track and the subsequent public-road test. Overall, the differences varied from mostly insignificant to low/medium in STOP SIGN and PUDO.